OpenStudy (elkiller33):

write -4 1/10 as decimal

ganeshie8 (ganeshie8):

\[- 4~ \dfrac{1}{10} ~~ =~~ - \left(4 + \dfrac{1}{10}\right) ~~=~~ -(4 + 0.1) ~~=~~ ?\]

OpenStudy (elkiller33):

what it equal? @ganeshie8

OpenStudy (vheah):

I am sure you can solve -(4 + 0.1)

OpenStudy (elkiller33):

i want the answer thooooooo :( @vheah

OpenStudy (vheah):

Yes and adding isn't hard to doooo. what is 4 + 0.1?

OpenStudy (onyxxelite):

-4.10